# Differentially Private 2D Human Pose Estimation

### v2

@@ 3.3. Gradient Subspace Identification
(6) $\tilde{g}_{i}=clip(\nabla l(w,z_{i}),C)$
@@
(7) $g=\frac{1}{B}\sum_{i\in B}\tilde{g}_{i}+\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I}),$
@@
3.4. Lightweight 2D-HPE: Integrating TinyViT Backbone with Coordinate Classification
For our 2D-HPE models, we adopt TinyViT (Wu et al., 2022) as the backbone. It is a small sized four-stage efficient hierarchical vision transformer which is well suited for resource-constrained vision tasks. The model adopts a multi-stage architecture where in the spatial resolution is progressively reduced and the feature representation expands. TinyViT follows a hybrid architectural design containing convolutional layers at the initial stages followed by self-attention mechanisms. Unlike standard ViT models, TinyViT employs a two-layer convolutional embedding. In the first stage of the network, it employs MBConv (Howard et al., 2019) blocks from MobileNetV2 to efficiently learn the low-level representation. The last three stages consists of transformer blocks hierarchically. Each stage consists of multi-head-self-attention (MHSA) layers, feed forward network (FFN) and $3$ x $3$ depthwise convolutions between the MHSA and FFN layers.
For keypoint localization, we augment the TinyViT backbone model with a coordinate classification output stage (Li et al., 2022). Given an input image $I\in\mathbb{R}^{C\times H\times W}$ and a ground truth keypoint $p_{i}=(x_{i},y_{i})$ for the $i^{\text{th}}$ joint, the continuous coordinates are quantized into discrete bins via a splitting factor $k\geq 1$. Formally, the quantized coordinates are computed as:
@@ 3.4. Lightweight 2D-HPE: Integrating TinyViT Backbone with Coordinate Classification
4. Experiments
4.1. Datasets and Evaluation Design
@@
In our experiments, we evaluated our framework on two widely used human pose datasets: MS COCO Keypoint Dataset (Lin et al., 2014) and MPII dataset (Andriluka et al., 2014). Our methodology assumes that the COCO dataset serves as a public dataset used for pre-training the network weights, while MPII functions as a private dataset on which we apply the differential privacy techniques.
@@
Specifically, our models are pretrained on the COCO train2017 set, which consists of approximately 118k images with around 140k annotated human instances, each with 17 joint annotations. The val2017 set consisting of around 5k images is used for validation. For evaluating the trade-off between utility and performance under various DP-SGD techniques we employ the MPII Human Pose Dataset consisting of 40k human instances, each labeled with 16 joint annotations. When transferring the model from COCO to MPII, we adjust for the keypoint discrepancy between datasets. We employ the Percentage of Correct Keypoints normalized by head (PCKh) (Andriluka et al., 2014) as an evaluation metric.
Our experimental framework explores three distinct DP training scenarios: Fine-Tuning with frozen backbone, Full Fine-Tuning and, Training from scratch. For the first scenario, we specifically freeze the first three stages of the backbone and finetune the fourth stage and all instances of layer norm(De et al., 2022). All models are trained under differential privacy constraints using DP-SGD with various clipping norms and privacy budgets ($\epsilon$). Each model undergoes training for a total of 25 epochs, as we empirically observed no significant performance improvements when extending training beyond this duration under DP constraints. Throughout all experiments, we maintain a fixed input resolution of $256\times 192$ pixels to ensure consistency across experiments and enable comparison with prior work.
For the privacy parameter settings, we use three gradient clipping norms $C=\{1.0,0.1,0.01\}$, with target privacy budgets of $\epsilon=\{0.2,0.4,0.6,0.8,1.0,3.0,5.0,7.0\}$, which range from tight privacy budgets to more relaxed. We adopt Renyi Differential Privacy (RDP) (Mironov, 2017) for privacy accounting with the privacy parameter $\delta=4e-5$.
4.2. Results
4.2.1. Baseline Results: Non-Private
Table 1 presents non-private baseline results from our model on the MPII dataset under three different training strategies: i) finetuning ii) finetuning from scratch and iii) training from scratch. These results demonstrate high pose estimation accuracy, representing best case performances in absence of DP noise.
Table 1. MPII Results: Non-Private Baselines for our HPE model on the MPII dataset
@@ 4.2.1. Baseline Results: Non-Private
4.2.2. Baseline Results with DP-SGD
Table 2 presents detailed performance results on the MPII dataset under the mentioned training strategies. The evaluation is carried out across multiple settings with different levels of privacy parameters $\epsilon$ and clipping thresholds $C$, and training strategies.
For standard finetuning with DP-SGD, we observe that lower clipping thresholds consistently result in better PCKh@0.5 scores across all joints. Specifically, with a clipping threshold of $C=0.01$, results substantially outperforms other thresholds, demonstrating that tighter gradient clipping significantly benefits model performance under strict privacy budgets. Notably, at $\epsilon=0.2$, mean PCKh@0.5 with $C=0.01$ reaches 63.85%, outperforming the 28.46% achieved with $C=0.1$ and the mere 5.94% with $C=1.0$. As expected, performance improves monotonically as $\epsilon$ increases across all joints and mean accuracy metrics. At a relaxed privacy constraint of $\epsilon=7.0$, PCKh@0.5 with the optimal clipping threshold ($C=0.01$) achieves 82.15%, highlighting the practical effectiveness of fine-tuning with a low clipping threshold for DP-SGD in pose estimation tasks.
In the scenario of fine-tuning from scratch, the overall performance noticeably deteriorates compared to standard fine-tuning. Under tight privacy constraints (e.g., $\epsilon=0.2$), performance is considerably limited across all clipping thresholds, with mean accuracy at $C=0.01$ dropping to 24.05%. Although, we observe similar patterns of improved accuracy values with larger $\epsilon$ and tighter clipping thresholds, the best performance in the settings ($\epsilon=7.0$, $C=0.01$) reaches only 55.66%, substantially lower than standard fine-tuning.
Training entirely from scratch presents the most significant challenge, with DP-SGD struggling to achieve competitive performance across all configurations. This demonstrates the inherent challenge of training deep neural networks privately without leveraging pretrained knowledge. Accuracy metrics across most body joints remain very low, rarely exceeding 20%, even at a weaker privacy constraints (say, $\epsilon=7.0$). The highest overall mean accuracy in this setting is a mere 17.37% at ($C=0.01$, $\epsilon=7.0$), underscoring the difficulty of privacy-preserving training with random initialization.
Overall, these results clearly indicate that fine-tuning pretrained models with small gradient clipping thresholds effectively reduces the performance degradation imposed by DP. Meanwhile, training from scratch remains remarkably challenging, even with careful DP-SGD configuration.
Table 2. MPII Results: DP-SGD.
@@ 4.2.2. Baseline Results with DP-SGD
$\epsilon=1.0$ 9.41 22.72 14.61 11.21 23.89 14.43 3.35 16.12 0.70
$\epsilon=3.0$ 9.79 19.92 16.82 11.43 22.61 15.19 4.09 15.99 0.72
$\epsilon=5.0$ 10.44 19.89 16.84 11.16 24.42 12.81 4.23 16.09 0.75
$\epsilon=7.0$ 15.04 22.06 17.27 11.82 23.63 14.27 6.92 17.37 0.78
4.2.3. Improved Results via Projection
Table 3 presents results on the MPII dataset for HPE using DP-SGD enhanced with the projection-based approach. Our experiments maintain fair training strategies and privacy parameters for direct comparison with baseline methods under DP-SGD in table 2.
In the fine-tuning scenario, we observe substantial improvements compared to approaches without projection. Even under stringent privacy budgets (low $\epsilon$ values), our method achieves high accuracy, particularly with lower clipping thresholds ($C=0.01$). For instance, at the strictest privacy budget ($\epsilon=0.2$) and the lowest clipping threshold ($C=0.01$), the mean accuracy reaches 78.48%, significantly outperforming both higher clipping thresholds (73.13% at $C=0.1$) and far exceeding the 10.34% at $C=1.0$. As privacy constraints become higher, performance improves consistently across all joints, with peak performance of 80.98% mean accuracy observed at $\epsilon=7.0$ and $C=0.01$. Notably, at moderate privacy budgets ($\epsilon=0.8$), we observe substantial accuracy gains for traditionally challenging joints to extract such as wrists and ankles.
In the setting of fine-tuning from scratch with DP projection, overall accuracies are markedly lower compared to freezing earlier layers. However, our proposed method still significantly outperforms corresponding results from the standard DP-SGD method without projection. Under strict privacy budgets (e.g., $\epsilon=0.2$), mean accuracy is limited, achieving a maximum of 53.54% at $C=0.01$. Performnace improves progressively with relaxed privacy constraints, reaching up to 69.81% accuracy at $\epsilon=7.0$ and $C=0.01$. The noticeable performance fluctuations at intermediate settings highlight the inherent sensitivity to both privacy budgets and clipping thresholds when fine-tuning from scratch.
For training entirely from scratch, limitations remain pronounced despite improvements over standard DP-SGD. Accuracies remain relatively low across all clipping thresholds and privacy budgets, though slightly outperforming standard DP-SGD without projection. Even at relaxed privacy budgets ($\epsilon=7.0$), the highest mean accuracy achieved is only 14.89% at $C=0.01$. These results highlight the fundamental challenge of fully private training without leveraging pretrained knowledge, even when employing gradient projection techniques.
Overall, our results presented in Table 3 strongly suggest that incorporating DP-SGD with projection significantly improves model performance, particularly when combined with strategic fine-tuning (freezing layers) and careful selection of gradient clipping thresholds. Among all evaluated approaches, fine-tuning with frozen initial layers at a low clipping threshold ($C=0.01$) emerges as the most effective strategy, achieving an excellent balance between model accuracy and privacy protection.
Table 3. MPII Results: DP-SGD with Projection.

### v3

@@ Introduction
$\tilde{g}_{i}=clip(\nabla l(w,z_{i}),C)$ (3)
@@
$g=\frac{1}{B}\sum_{i\in B}\tilde{g}_{i}+\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I}),$ (4)
@@
Feature Projective DP-SGD for HPE
To enhance the model utility in our 2D HPE task, we extend the standard DP-SGD framework using Feature Differential Privacy (FDP). FDP exploits the transformation of the training image into private and public variants, selectively applying differential privacy only to sensitive features while freely utilizing non-sensitive(public) information (Mahloujifar et al. 2025). Formally, let each sample be $x_{i}\in S_{data}$(a raw training image with keypoint labels), and let $\psi:S\rightarrow\mathcal{F}$ be a public feature map such that $\psi(x_{i})$ is the public variant of the raw image $x_{i}$ and let $f:[0,1]\rightarrow[0,1]$ be a trade-off function. A randomized mechanism $\mathcal{M}$ satisfies $f$-FDP with respect to $\psi$ if, for any two datasets $d,d^{\prime}$ differing in exactly one image-label pair $x_{i}\neq x_{i}^{\prime}$ but having identical public representations $(\psi(x_{i})=\psi(x_{i}^{\prime}))$ and for all subsets $S$ for range of $\mathcal{M}$:
$\mathbb{P}[\mathcal{M}(d)\in S]\leq 1-f(\mathbb{P}[\mathcal{M}(d^{\prime})\in S])$ (6)
Then, we say the mechanism is $(\epsilon,\delta)$-DP with respect to $\psi$ iff it is $f$-FDP for $f(x)=1-\delta-e^{\epsilon}x$. Motivated by this definition, the FDP-SGD method, explicitly distinguishes between public and private (raw) images to improve pose estimation accuracy under the same privacy budget as standard DP. Specifically, for each image-keypoint pair, we define a public loss $l_{pub}(w,\psi(x))$ which captures the coarse pose estimation based on the definition of $\psi$. The private loss $l_{priv}(w,x)$ captures the sensitive, fine-grained details of the human that requires privacy protection. Then the overall loss can be given as:
$l(w,x)=l_{priv}(w,x)+l_{pub}(w,\psi(x))$ (7)
Training:
At every iteration, FDP-SGD updates parameters using two separate batches drawn independently from $S_{data}$. On the public batch $B_{pub}$ we compute the gradient as:
$g_{pub}=\frac{1}{|B_{pub}|}\sum_{x\in B_{pub}}\nabla l_{pub}(w,\psi(x))$ (8)
Similarly, on the private batch $B_{priv}$, we compute and clip the gradient of the private loss to the clipping norm $C$ as $\tilde{g}$, then aggregate and add gaussian noise:
$g_{priv}=\frac{1}{|B_{priv}|}\sum_{x\in B_{priv}}\tilde{g}+\mathcal{N}(0,\sigma^{2}C^{2}I)$ (9)
The parameters are updated at iteration $t$ by combining both the public and private gradients as:
$g=g_{priv}+g_{pub}$ (10)
$w_{t}=w_{t-1}-\eta_{t}g_{t}$ (11)
where $\eta_{t}$ denotes the learning rate. By applying noise only to the private gradient component, FDP-SGD significantly improves the accuracy of human pose estimation compared to vanilla DP-SGD under identical $(\epsilon,\delta)$ privacy constraints.
To further enhance utility, we integrate FDP-SGD with our previously described subspace projection technique, creating Feature-Projective DP. This combined approach focuses parameter updates on the most informative gradient directions while maintaining the privacy benefits of feature decomposition. The complete algorithmic details of this integrated approach are provided in Algorithm 1.
Algorithm 1 Feature Projective DP-SGD
0: Full dataset $\mathcal{S}_{\rm data}=\{x_{1},\dots,x_{n}\}$, split into public subset $\mathcal{S}_{\rm pub}\subset\mathcal{S}_{\rm data}$ (size $m$) and private remainder $\mathcal{S}_{\rm priv}=\mathcal{S}_{\rm data}\setminus\mathcal{S}_{\rm pub}$, public feature map $\psi$,combined loss $\ell(w,z)$ with public and private losses $l_{pub},l_{priv}$, clip norm $C$, noise std. $\sigma$, subspace dim $k$,batch size $B$, iterationss $T$, learning rate $\{\eta_{t}\}$.
1: Initialize model parameters $w_{0}\in\mathbb{R}^{p}$.
2:for $t=1,\dots,T$ do
3:  (1) Subspace identification on $\mathcal{S}_{\rm pub}$:
4:  Compute
@@
5:  Compute the top-$k$ eigenvectors of $M_{t}$.
6:  Form the subspace basis $V_{t}\in\mathbb{R}^{p\times k}$.
7:  Compute the projector $\hat{V_{t}}\,\hat{V_{t}}^{\top}\in\mathbb{R}^{p\times p}$.
8:  (2) Compute public and private feature gradient:
9:  Sample public batch $B_{pub}^{t}\subset{\psi(x):x\in S_{priv}}$
10:  Compute
$g_{pub}^{t}=\frac{1}{|B_{pub}^{t}|}\sum_{x\in B_{pub}^{t}}\nabla l_{pub}(w_{t-1},\psi(x))$
11:  Sample private batch $B_{priv}^{t}\subset S_{priv}$
12:  Compute the clipped gradient $\tilde{g_{t}}$, aggregate and add Gaussian noise
$g_{priv}^{t}=\frac{1}{|B_{priv}^{t}|}\sum_{x\in B_{priv}^{t}}\tilde{g_{t}}+\mathcal{N}(0,\sigma^{2}C^{2}I)$
[…]
14:  Merge Public and Private projected feature gradients
$g_{t}=g_{pub}^{t}+g_{proj}^{t}$
[…]
16:end for
17:return $w_{T}$.
Experiments
Dataset and Implementation Details
@@
In our experiments, we evaluated our framework on two widely used human pose datasets: MS COCO Keypoint Dataset (Lin et al. 2014) and MPII dataset (Andriluka et al. 2014). Our methodology assumes that the COCO dataset serves as a public dataset used for pre-training the network weights, while MPII functions as a private dataset on which we apply the differential privacy techniques. The MS COCO Keypoint Dataset is employed for pretraining purposes.
@@
Specifically, our models are pretrained on the COCO train2017 set, which consists of approximately 118k images with around 140k annotated human instances, each with 17 joint annotations. The val2017 set consisting of around 5k images is used for validation. For evaluating the trade-off between utility and performance under various DP-SGD techniques we employ the MPII Human Pose Dataset consisting of 40k human instances, each labeled with 16 joint annotations. When transferring the model from COCO to MPII, we adjust for the keypoint discrepancy between datasets. We employ the Percentage of Correct Keypoints normalized by head (PCKh) (Andriluka et al. 2014) as an evaluation metric. To further assess the generalization ability of our pose estimation models under domain shift and visual diversity, we conduct additional experiments on the Human-Art dataset (Ju et al. 2023). Human-Art is a recently introduced, large-scale human-centric dataset specifically designed to bridge the gap between natural and artificial visual domains. It contains 50,000 high-quality images with over 123,000 person instances drawn from 20 diverse scenarios, covering both natural scenes (e.g., cosplay, drama, dance) and a wide spectrum of artistic representations (e.g., oil paintings, sculptures, digital art, watercolor, and murals).
Compared to conventional datasets like MPII or COCO, Human-Art presents significantly greater challenges for pose estimation. This is primarily due to the presence of stylized or abstract depictions of human figures, exaggerated or distorted body proportions, occlusions, artistic textures, and unconventional pose. The results on HumanART are presented in supplementary.
Our experimental framework explores three distinct DP training scenarios: Fine-Tuning with frozen backbone, Full Fine-Tuning and, Training from scratch. For the first scenario, we specifically freeze the first three stages of the backbone and finetune the fourth stage and all instances of layer norm(De et al. 2022). To generate the public feature map, we employ Gaussian blur which effectively suppresses facial and body structure details. Details on datasets, training and privacy related parameters are provided in the supplementary.
Figure 2: Comparison of PCKh@0.5 across private and non-private methodologies under different training strategies with varied privacy budget ($\epsilon$) and clipping thresholds ($C$).
Experimental Results and Analysis
For comprehensive comparison, all experimental results across training strategies, clipping thresholds and privacy budgets are visualized in Figure 2.
Non-Private Baseline Results:
Table 1 presents baseline pose estimation performance of our model on the MPII dataset under three training strategies: (i) finetuning from a COCO-pretrained model, (ii) finetuning from scratch (initialization with COCO pretrained weights and all layers are trained), and (iii) training from scratch (random initialization). Additionally, we report results from using only public features (blurred images) under the same strategies to provide context for evaluating privacy-utility trade-offs.
Table 1: MPII Results: Non-Private Baselines for our HPE model on the MPII dataset
@@
Finetuning on Public features 94.30 93.95 83.21 75.19 83.53 76.86 72.76 83.61 20.81
Finetuning from scratch on Public features 88.71 84.32 69.71 54.37 70.24 61.67 55.12 70.32 11.36
Training from scratch on Public features 15.31 20.01 17.19 12.59 25.04 16.90 10.91 17.99 0.78
As expected, the finetuning strategy achieves the highest mean accuracy of 89.36% followed by finetuning from scratch (88.28%) and training from scratch (76.89%), which is to be expected. These non-private baselines establish upper bound performance references for evaluating differential privacy impact. When the model relies only on public features(gaussian blurred images), performance reduces significantly. While finetuning on public features maintain reasonable accuracy, the other training strategies yield substantially compromised results, confirming that fine-grained visual details in raw images are critical for accurate pose estimation.
Figure 3: Depiction of qualitative results on DP-SGD, Projection DP-SGD and Feature Projection DP-SGD. We specifically show results on Finetuning with $C=0.1$ at various privacy budgets.
DP-SGD Baseline Results
Figure 2 presents the PCKh@0.5 results on MPII dataset under the aforementioned training strategies using DP-SGD. Experiments were conducted across multiple settings with varying privacy parameters ($\epsilon\in\{0.2,0.4,0.6,0.8\}$) and clipping thresholds ($C\in\{0.01,0.1,1.0\}$). For standard finetuning with DP-SGD, lower clipping thresholds consistently yield better pose estimation results across different privacy levels. Specifically, at $C=0.01$, the model achieves substantially higher accuracy of 63.85% mean PCKh@0.5 at the tightest privacy loss($\epsilon=0.2$) compared to $C=0.1$ (28.46%) and $C=1.0$ (5.94%). This is indeed because of the fact that the effective noise magnitude grows linearly with the $C$ thus our results confirm this.
Notably, finetuning the COCO-pretrained TinyViT backbone significantly mitigates the DP induced performance degradation compared to training from scratch or finetuning from scratch (Yu et al. 2021). This indicates that pretrained human pose based feature representations provide robust feature priors that enable DP-SGD to adapt effectively to private pose datasets, while maintaining resilience to noise corruption.
Performance Analysis of Subspace Projection
We maintain identical training strategies and privacy parameters to ensure direct comparison with both non-private and DP-SGD baseline methods. Our subspace projection approach demonstrates substantial performance improvements across multiple configurations. At the most restrictive clipping threshold ($C=0.01$), projection yields significant gains from 63.85% to 78.48% at $\epsilon=0.2$ and from 78.17% to 80.63% at $\epsilon=0.8$. This enhancement occurs because, while Gauissian noise is injected uniformly across all gradient components, only a subset of directions carry meaningful pose-relevant information. By projecting onto the learned subspace, we effectively discard the noise in irrelevant directions, thereby improving signal-to-noise ratio and preserving essential pose estimation features. At $C=0.1$, the projection approach consistently outperforms baseline DP-SGD. Finetuning increases accuracy from 73.13% to 77.41%, while training from scratch improves from 9.80% to 13.05%. However, for the finetuning from scratch strategy, the curve plateaus slightly below regular DP-SGD. We attribute this phenomenon to the interaction between injected gaussian noise and subspace reconstruction error(Zhou, Wu, and Banerjee 2020). At the largest clipping threshold ($C=1.0$), we observe non-monotonic patterns. Under this condition, the raw gradients become dominated by noise, leading to unstable parameter updates and local dips in accuracy.
Performance analysis of FDP and Feature-Projective DP
Feature DP consistently outperforms vanilla DP-SGD across all experimental configurations. Under finetuning with $C=0.01$, FDP achieves substantial improvements, from 63.85% to 75.46% at $\epsilon=0.2$ (11.61% gain) and from 78.17% to 80.40% at $\epsilon=0.8$ (2.23% gain). This is consistently observed across all training strategies and clipping values. Integrating FDP with subspace projection results in the highest accuracy across all experimental settings. Even under the most challenging conditions with stringent clipping of $C=1.0$, where standard DP-SGD achieves only 12.53%, Feature-projective DP attains 71.66%, representing a six fold relative gain.
The largest improvements occur when training from scratch. With $C=0.1$ and $\epsilon=0.8$, vanilla DP-SGD achieves merely 6.85% accuracy, while FDP alone attains 11.22%. However, the combined Feature-Projective DP approach achieves 33.48%. This demonstrates that combining both techniques boosts utility drastically especially in large noise induced scenarios, where neither alone suffices to recover strong pose features from corrupted gradients. Figure 4 depicts few qualitative results across different privacy strategies along with ground truth.
For completeness, detailed per-joint performance metrics, full qualitative results and additional experimental results are provided in the Supplementary Material.
Conclusion
Our work presents the first differentially private (DP) approach to 2D human pose estimation (HPE), addressing critical privacy concerns while maintaining utility. Our results clearly establish that the synergistic combination of feature-level privacy and subspace projection dramatically enhances utility across all settings. Importantly, our proposed Feature-Projective DP 2D-HPE approach achieved up to 82.61% mean PCKh@0.5 at $\epsilon=0.8$, significantly narrowing the gap to non-private performance while exhibiting strong formal privacy guarantees. A key advantage of this method is that it requires no manual curation of private features, as we automatically protect the complete raw image, ensuring privacy preservation for both individuals and their personal spatial environments.
Appendix A Implementation Details
Our models are pretrained on the COCO train2017 set, which consists of approximately 118k images with around 140k annotated human instances, each with 17 joint annotations. The val2017 set consisting of around 5k images is used for validation. For evaluating the trade-off between utility and performance under various DP-SGD techniques we employ the MPII Human Pose Dataset consisting of 40k human instances, each labeled with 16 joint annotations. When transferring the model from COCO to MPII, we adjust for the keypoint discrepancy between datasets. We employ the Percentage of Correct Keypoints normalized by head (PCKh) (Andriluka et al. 2014) as an evaluation metric.
All models are trained under differential privacy constraints using DP-SGD with various clipping norms and privacy budgets ($\epsilon$). Each model undergoes training for a total of 25 epochs, as we empirically observed no significant performance improvements when extending training beyond this duration under DP constraints. Throughout all experiments, we maintain a fixed input resolution of $256\times 192$ pixels to ensure consistency across experiments and enable comparison with prior work.
For the privacy parameter settings, we use three gradient clipping norms $C=\{1.0,0.1,0.01\}$, with target privacy budgets of $\epsilon=\{0.2,0.4,0.6,0.8\}$. We adopt Renyi Differential Privacy (RDP) (Mironov 2017) for privacy accounting with the privacy parameter $\delta=4e-5$.
To generate the public feature map, we employ Gaussian blur as $\psi$ with a kernel size of $(25,25)$ and standard deviation $\sigma_{X}=10$, which effectively suppresses facial and body structure details. We deliberately blur the entire image rather than selectively masking human regions, as this approach provides comprehensive privacy protection by obscuring not only human identities but also contextual environmental details.
We chose to blur the complete image owing to the technicalities of the dataset as MPII has multiple instances of humans from similar image.
Appendix B Per Joint PCKh@0.5 on MPII and HumanART
Tables below report per-joint PCKh@0.5 scores on MPII under various DP-SGD settings. Smaller clipping norms (C) and higher privacy budgets ($\epsilon$) consistently improve pose accuracy. While we provide detailed results on the MPII dataset across various DP-SGD settings, including training from scratch, fine-tuning, and projection-based setups under different clipping norms and privacy budgets, we only report a subset of results for the Human-Art dataset. This is because certain settings, especially those involving large clipping norms or training without pretraining, result in extremely low utility. In many of these cases, the models fail to converge or produce meaningful predictions, with PCKh scores often dropping below 5%. Since these settings do not offer useful insights for practical analysis, we exclude them from the main tables and focus on configurations that yield more stable and interpretable performance
Table 2: MPII Results: DP-SGD.
@@ Appendix A Implementation Details
Table 3: MPII Results: DP-SGD with Projection.